OpenStudy (anonymous):

Find the solution to each system by the substitution method for these, 10x3y=8 and 2x+y=2.......also have so how to check your answers......so far i have 10x+3y=8, 10x=3y+8,2(3y+8)+y=8 6y+16+y=8, 7y+16=8, 7y=16+2 7y=18, y=2.5

OpenStudy (anonymous):

x = 1/2 , y = 1

OpenStudy (anonymous):

ok so what did i do wrong then

OpenStudy (anonymous):

10x = 3y +8 is wrong..........

OpenStudy (anonymous):

ok i dont get how that is wrong

OpenStudy (anonymous):

negative sign

OpenStudy (anonymous):

so would it be 10x=3y-8

OpenStudy (anonymous):

10x+3y=8 then it become 10x = -3y+8 right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you want me explain all?

OpenStudy (anonymous):

preferrably if you would

OpenStudy (anonymous):

im not very good but i will try my best

OpenStudy (anonymous):

10x+3y=8 and 2x+y=2 take 2x+y=2 to be y =2-2x then insert y =2-2x into 10x+3y=8 10x+3(2-2x)=8 10x+6-6x=8 4x+6=8 4x=8-6 4x=2 x=2/4 x=1/2 then insert x=1/2 into y =2-2x y=2-2(1/2) y=2-1 y=1

OpenStudy (anonymous):

ok thank you!